The vapour pressure of pure water is 760 mm at 25 degree Celsius the the vapour pressure of solution containing 1 m solution of glucose will be

Grade:12

2 Answers

Arun
4 years ago
Dear student
 
Molality 1 m
 
1 * 10^-3 * 18 = 760 -p /760
 
p = 746.3w mm Hg
 
Hope it helps
 
Regards
Arun (askIITians forum expert)
SWAYAM PRAKASH MALIK
2 years ago
According to the question ,
Molality = 1 m ,
Now , 760-p/p = 1 * 10^-3 * 18 ,
=> p = 746.3 mm Hg . 
Hope it helps.